# HD3C: Efficient Medical Data Classification for Embedded Devices

### v1

@@ Abstract
Energy-efficient medical data classification is essential for modern disease screening, particularly in home and field healthcare where embedded devices are prevalent. While deep learning models achieve state-of-the-art accuracy, their substantial energy consumption and reliance on GPUs limit deployment on such platforms. We present Hyperdimensional Computing with Class-Wise Clustering (HD3C), a lightweight classification framework designed for low-power environments. HD3C encodes data into high-dimensional hypervectors, aggregates them into multiple cluster-specific prototypes, and performs classification through similarity search in hyperspace. We evaluate HD3C across three medical classification tasks; on heart sound classification, HD3C is $350\times$ more energy-efficient than Bayesian ResNet with less than 1% accuracy difference. Moreover, HD3C demonstrates exceptional robustness to noise, limited training data, and hardware error, supported by both theoretical analysis and empirical results, highlighting its potential for reliable deployment in real-world settings. Code is available here.
@@ 1 Introduction
To address these challenges, we present Hyperdimensional Computing with Class-Wise Clustering (HD3C), a lightweight classification framework that extends the standard Hyperdimensional Computing (HDC). HD3C encodes each sample into a high-dimensional sample hypervector (Sample-HV), aggregates them into a compact set of cluster prototypes (Cluster-HVs), and performs classification by selecting the Cluster-HV with the highest similarity.
@@
    We propose HD3C, a lightweight classification framework extended from Hyperdimensional Computing (HDC). It achieves significantly higher accuracy than standard HDC and other efficient models while preserving energy efficiency.
@@
    We demonstrate that HD3C achieves substantially lower energy consumption than deep learning baselines across three medical tasks. On heart sound classification, it is $350\times$ more energy-efficient than Bayesian ResNet while maintaining comparable accuracy.
@@
    We provide both theoretical and empirical evidence of HD3C’s robustness on input noise and hardware error. To the best of our knowledge, our theorems on binary hypervectors, although some are widely recognized, have not been formally proven in prior HDC research.
This paper is organized as follows. Section 2 reviews the fundamentals of hyperdimensional computing. Section 3 details the proposed HD3C framework. Section 4 presents experimental results and sensitivity analysis, and Section 5 concludes the paper.
@@ 2 Hyperdimensional Computing (HDC)
3 Medical Data Classification through HD3C
We propose Hyperdimensional Computing with Class-Wise Clustering (HD3C), an extension of the HDC framework that improves accuracy while preserving energy efficiency. Algorithm 1 and Figure 3 illustrate the overall HD3C pipeline. In this work, medical samples are first encoded into sample hypervectors (Section 3.1). These sample hypervectors are then clustered into a compact set of cluster prototypes, and classification is performed by similarity search against these prototypes (Section 3.2). An optional retraining stage may further enhance accuracy (Section 3.3). Formal definitions, lemma, and theorems supporting this framework are provided in Appendix A.
@@ 3.1 Encode Sample into Hypervector
HD3C requires numeric features as input. For heart sound classification, we extract $d=720$ features using Mel-frequency Cepstral Coefficients (MFCC) (Davis & Mermelstein, 1980) and Discrete Wavelet Transform (DWT) (Mallat, 1989), two widely used frequency-domain representations in audio analysis. For breast cancer classification, we use $d=30$ features from fine-needle aspirate (FNA) breast mass images, and for EMG classification, we use $d=8$ features.
@@
Figure 3: Medical data classification through HD3C: Training samples are encoded into sample hypervectors (Sample-HVs) and aggregated into a compact set of cluster prototypes (Cluster-HVs); new samples are classified by selecting the Cluster-HV with highest similarity. The figure illustrates a binary classification example, though HD3C is not limited to binary tasks.
@@ 3.2 Class-Wise Hyperspace Clustering
To address this, HD3C introduces class-wise hyperspace clustering. The clustering process is inspired by K-means, but performed in hyperspace on Sample-HVs and applied independently within each class. Unlike K-means which computes arithmetic means, HD3C uses the bundling operation (Definition 2) to form cluster prototypes (Cluster-HVs). Leveraging hypervectors rather than raw numeric features provides the robustness formalized in Theorem 1.
@@
Algorithm 1 HD3C: Hyperdimensional Computing with Class-Wise Clustering
@@
This bundling and reassigning process is repeated for $T$ iterations until convergence, yielding the final set of Cluster-HVs $\{{C}_{j}^{k}\}^{K}_{k=1}$ for class $j$. After clustering all $J$ classes, HD3C may classify an unseen new sample by encoding it as Sample-HV $S^{\mathrm{new}}$ and assigning it to the Cluster-HV with the highest similarity (i.e., lowest Hamming distance):
@@ 3.3 Retrain Cluster Prototypes
To further improve accuracy, HD3C optionally applies a retraining procedure that adjusts Cluster-HVs based on misclassified training samples. The retraining stage operates exclusively on the training set with no reference to test data.
@@
To correct this, HD3C performs two adjustments: (1) Subtracts $S^{\mathrm{err}}$ from the incorrect Cluster-HV $C_{j_{2}}^{k_{2}}$ and (2) Re-bundles $S^{\mathrm{err}}$ onto the correct Cluster-HV $C_{j_{1}}^{k_{1}}$. We generalize this operation across all misclassified training samples. Let $\mathcal{E}_{j}^{\mathrm{k,out}}$ denote Sample-HVs outside cluster $\mathfrak{C}_{j}^{k}$ that incorrectly match closest to $C_{j}^{k}$; while $\mathcal{E}_{j}^{\mathrm{k,in}}$ denote the Sample-HVs within $\mathfrak{C}^{j}_{k}$ but are closer to a Cluster-HV of a different cluster. The retrained Cluster-HV representing cluster $\mathfrak{C}^{j}_{k}$ is computed as
@@ 4.2 Experiment Setup
Platform: Large deep learning baselines were trained and evaluated on an NVIDIA GeForce RTX 4090 GPU (24GB VRAM). HD3C, other one-shot models, and efficient neural networks were trained and evaluated on an Intel Xeon Gold 6430 CPU (120GB RAM). All experiments were conducted on Ubuntu 22.04.1 LTS (Linux kernel 5.15.0-97-generic).
@@ 4.3 Comparison with Baseline Models
Table 1 summarizes the performance of HD3C against baseline models across three medical datasets. HD3C achieves substantially higher energy efficiency than deep neural networks and outperforms other efficient methods in terms of accuracy. For instance, on the heart sounds classification task, HD3C is $580\times$ more energy-efficient in training and $350\times$ in inference compared to the best-performing Bayesian ResNet (H. Krones et al., 2022), while also delivering a 10% accuracy improvement over standard HDC.
Table 1: Performance and Energy Comparison of HD3C and Baseline Models
@@
HD3C (Ours) 91.592 ± 2.927 2.813 ±   0.205 0.422 ± 0.029
4.4 Impact of HD3C Hyperparameters
As discussed in Section 3, a sufficiently large dimensionality $D$ is critical for HD3C to maintain robustness against input noise (Theorem 1) and to ensure that Cluster-HVs accurately capture the aggregate features of their clusters (Theorem 2). Likewise, an adequate number of clusters $K$ are essential for forming stable hyperspace clusters. A few retrain epochs can further fine-tune Cluster-HVs to better align with the Sample-HVs assigned to each cluster.
However, excessively large values of these hyperparameters introduce unnecessary computational overhead, reducing the energy efficiency HD3C is designed to achieve. Our experiments also indicate that allocating too many clusters or performing excessive retraining can lead to overfitting. Figure 4 (a-c) illustrates HD3C’s performance on heart sound classification across different hyperparameter settings.
@@ 4.5.1 Resilience to Input Noise
Real-world medical signals often contain persistent environmental noise, which can impair classification performance (Clifford et al., 2016). The inherent robustness of HD3C’s encoding mechanism, as supported by Theorem 1, allows it to maintain high accuracy even with noisy inputs. As shown in Figure 4(d), HD3C experiences only 1.39% drop in accuracy under 15% input noise, demonstrating its reliability in noisy settings.
@@ 4.5.2 Resilience to Limited Training Data
Many medical datasets are limited in size due to privacy constraints. As shown in Figure 4(e), HD3C remains robust under reduced training data, with only a 1.78% drop in accuracy when trained on 40% of the PhysioNet 2016 dataset.
@@ 4.5.3 Resilience to Hardware Errors
To evaluate this robustness in HD3C, we conducted a perturbation analysis by randomly flipping elements (+1 to -1 and vice versa) in all stored Cluster-HVs to simulate hardware instability. Theorem 3 demonstrates that, with sufficient dimensionality, flipping up to 50% of the elements has minimal impact on classification accuracy (Proof 3). The theorem is supported empirically: as shown in Figure 4(f), with D=10, [^0] D=10{,}000, flipping 20% of elements results in only a 2.84% drop in accuracy.
Figure 4: HD3C sensitivity to hyperparameters, input noise, limited training data, and hardware errors on PhysioNet 2016.
@@ 4.6 Conceptual Hardware Framework
Figure 5: Conceptual hardware framework for HD3C.
The current implementation of HD3C is in Python, which does not fully exploit the hardware-level efficiency of binary hypervectors enabled by single-bit operations and parallel computing. Prior work on HDC hardware has shown that fundamental operations such as binding and bundling can be implemented with extremely lightweight digital logic: binding reduces to bitwise XOR, bundling to majority voting circuits, and similarity search to parallelized Hamming distance computation (Imani et al., 2017). These properties allow HDC systems to operate with low energy consumption and high throughput compared to conventional floating-point ML models.
Building on these insights, we outline a conceptual hardware framework specifically tailored for HD3C (Figure 5). To validate its feasibility, we prototyped the framework on a Xilinx FPGA platform and are able to achieve the accuracy reported in Table 1. We plan to further refine and evaluate this hardware-oriented design in future work.
@@ 5 Conclusion
This paper introduces HD3C, an energy-efficient classification framework extended from Hyperdimensional Computing (HDC). HD3C demonstrates significant advantage in medical data classification: on heart sound classification, it is $350\times$ more efficient than Bayesian ResNet, and provide $10\%$ accuracy improvement over standard HDC.
We evaluated HD3C’s robustness both theoretically and empirically. The model maintains high accuracy under challenging conditions: showing only a 1.39% drop under 15% input noise, 1.78% when trained on just 40% of the dataset, and 2.84% when 20% of its learned parameters are corrupted. These results highlight HD3C’s reliability on real-world medical applications.
From a societal perspective, HD3C can help expand access to medical screening in underserved settings by enabling assessments on low-cost, GPU-free devices. However, overreliance on the screening tool without clinical oversight could lead to misdiagnoses; therefore, the system should be designed to recommend follow-up with a medical professional when abnormalities are detected.
@@
We are committed to ensuring the reproducibility of our results. To this end, we have released the complete HD3C codebase along with a detailed installation guide in the Supplementary Materials. The repository contains all datasets used in our experiments as well as the exact hyperparameter configurations adopted for training and evaluation. This enables researchers to fully replicate our experiments and verify the reported results.

### v2

# HDC-X: Efficient Medical Data Classification for Embedded Devices

@@ Abstract
Energy-efficient medical data classification is essential for modern disease screening, particularly in home and field healthcare where embedded devices are prevalent. While deep learning models achieve state-of-the-art accuracy, their substantial energy consumption and reliance on GPUs limit deployment on such platforms. We present HDC-X, a lightweight classification framework designed for low-power devices. HDC-X encodes data into high-dimensional hypervectors, aggregates them into multiple cluster-specific prototypes, and performs classification through similarity search in hyperspace. We evaluate HDC-X across three medical classification tasks; on heart sound classification, HDC-X is $350\times$ more energy-efficient than Bayesian ResNet with less than 1% accuracy difference. Moreover, HDC-X demonstrates exceptional robustness to noise, limited training data, and hardware error, supported by both theoretical analysis and empirical results, highlighting its potential for reliable deployment in real-world settings. Code is available here.
@@ 1 Introduction
To address this, we present Hyperdimensional Computing - Next Generation (HDC-X), a lightweight classification framework that extends standard Hyperdimensional Computing (HDC) while significantly improving its performance. HDC-X encodes each sample into a high-dimensional sample hypervector (Sample-HV), aggregates them into a compact set of cluster prototypes (Cluster-HVs), and performs classification by selecting the Cluster-HV with the highest similarity.
@@
    We propose HDC-X, a lightweight classification framework extended from Hyperdimensional Computing (HDC). It achieves significantly higher accuracy than standard HDC and other efficient models while preserving energy efficiency.
@@
    We demonstrate that HDC-X consumes far less energy than deep learning baselines across three medical tasks. On heart sound classification, it is $350\times$ more energy-efficient than Bayesian ResNet while maintaining comparable accuracy.
@@
    We provide both theoretical and empirical evidence of HDC-X’s robustness to input noise and hardware error. To the best of our knowledge, our theorems on binary hypervectors, although some are widely recognized, have not been formally proven in prior HDC research.
This paper is organized as follows. Section 2 reviews the fundamentals of hyperdimensional computing. Section 3 details the proposed HDC-X framework. Section 4 presents experimental results and sensitivity analysis, and Section 5 concludes the paper.
@@ 2 Hyperdimensional Computing (HDC)
3 Medical Data Classification through HDC-X
We propose Hyperdimensional Computing Next Generation (HDC-X), an extension of the HDC framework that significantly improves accuracy while preserving energy efficiency. Algorithm 1 and Figure 3 illustrate the overall HDC-X pipeline. In this work, medical samples are first encoded into sample hypervectors (Section 3.1). These sample hypervectors are then clustered into a compact set of cluster prototypes, and classification is performed by similarity search against these prototypes (Section 3.2). An optional retraining stage may further enhance accuracy (Section 3.3). Formal definitions, lemma, and theorems supporting this framework are provided in Appendix A.
@@ 3.1 Encode Sample into Hypervector
HDC-X requires numeric features as input. For heart sound classification, we extract $d=720$ features using Mel-frequency Cepstral Coefficients (MFCC) (Davis & Mermelstein, 1980) and Discrete Wavelet Transform (DWT) (Mallat, 1989), two widely used frequency-domain representations in audio analysis. For breast cancer classification, we use $d=30$ features from fine-needle aspirate (FNA) breast mass images, and for EMG classification, we use $d=8$ features.
@@
Figure 3: Medical data classification through HDC-X: Training samples are encoded into sample hypervectors (Sample-HVs) and aggregated into a compact set of cluster prototypes (Cluster-HVs); new samples are classified by selecting the Cluster-HV with highest similarity. The figure illustrates a binary classification example, though HDC-X is not limited to binary tasks.
@@ 3.2 Class-Wise Hyperspace Clustering
To address this, HDC-X introduces class-wise hyperspace clustering. The clustering process is inspired by K-means, but performed in hyperspace on Sample-HVs and applied independently within each class. Unlike K-means which computes arithmetic means, HDC-X uses the bundling operation (Definition 2) to form cluster prototypes (Cluster-HVs). Leveraging hypervectors rather than raw numeric features provides the robustness formalized in Theorem 1.
@@
Algorithm 1 HDC-X: Hyperdimensional Computing - Next Generation (Without Retraining)
@@
This bundling and reassigning process is repeated for $T$ iterations until convergence, yielding the final set of Cluster-HVs $\{{C}_{j}^{k}\}^{K}_{k=1}$ for class $j$. After clustering all $J$ classes, HDC-X may classify an unseen new sample by encoding it as Sample-HV $S^{\mathrm{new}}$ and assigning it to the Cluster-HV with the highest similarity (i.e., lowest Hamming distance):
@@ 3.3 Retrain Cluster Prototypes
To further improve accuracy, HDC-X optionally applies a retraining procedure that adjusts Cluster-HVs based on misclassified training samples. The retraining stage operates exclusively on the training set with no reference to test data.
@@
To correct this, HDC-X performs two adjustments: (1) Subtracts $S^{\mathrm{err}}$ from the incorrect Cluster-HV $C_{j_{2}}^{k_{2}}$ and (2) Re-bundles $S^{\mathrm{err}}$ onto the correct Cluster-HV $C_{j_{1}}^{k_{1}}$. We generalize this operation across all misclassified training samples. Let $\mathcal{E}_{j}^{\mathrm{k,out}}$ denote Sample-HVs outside cluster $\mathfrak{C}_{j}^{k}$ that incorrectly match closest to $C_{j}^{k}$; while $\mathcal{E}_{j}^{\mathrm{k,in}}$ denote the Sample-HVs within $\mathfrak{C}^{j}_{k}$ but are closer to a Cluster-HV of a different cluster. The retrained Cluster-HV representing cluster $\mathfrak{C}^{j}_{k}$ is computed as
@@ 4.2 Experiment Setup
Platform: Large deep learning baselines were trained and evaluated on an NVIDIA GeForce RTX 4090 GPU (24GB VRAM). HDC-X, other one-shot models, and efficient neural networks were trained and evaluated on an Intel Xeon Gold 6430 CPU (120GB RAM). All experiments were conducted on Ubuntu 22.04.1 LTS (Linux kernel 5.15.0-97-generic).
@@ 4.3 Comparison with Baseline Models
Table 1 summarizes the performance of HDC-X against baseline models across three medical datasets. HDC-X achieves substantially higher energy efficiency than deep neural networks and outperforms other efficient methods in terms of accuracy. For instance, on the heart sounds classification task, HDC-X is $580\times$ more energy-efficient in training and $350\times$ in inference compared to the best-performing Bayesian ResNet (H. Krones et al., 2022), while also delivering a 10% accuracy improvement over standard HDC.
Table 1: Performance and Energy Comparison of HDC-X and Baseline Models
@@
HDC-X (Ours) 91.592 ± 2.927 2.813 ±   0.205 0.422 ± 0.029
4.4 Impact of HDC-X Hyperparameters
As discussed in Section 3, a sufficiently large dimensionality $D$ is critical for HDC-X to maintain robustness against input noise (Theorem 1) and to ensure that Cluster-HVs accurately capture the aggregate features of their clusters (Theorem 2). Likewise, an adequate number of clusters $K$ are essential for forming stable hyperspace clusters. A few retrain epochs can further fine-tune Cluster-HVs to better align with the Sample-HVs assigned to each cluster.
However, excessively large values of these hyperparameters introduce unnecessary computational overhead, reducing the energy efficiency HDC-X is designed to achieve. Our experiments also indicate that allocating too many clusters or performing excessive retraining can lead to overfitting. Figure 4 (a-c) illustrates HDC-X’s performance on heart sound classification across different hyperparameter settings.
@@ 4.5.1 Resilience to Input Noise
Real-world medical signals often contain persistent environmental noise, which can impair classification performance (Clifford et al., 2016). The inherent robustness of HDC-X’s encoding mechanism, as supported by Theorem 1, allows it to maintain high accuracy even with noisy inputs. As shown in Figure 4(d), HDC-X experiences only 1.39% drop in accuracy under 15% input noise, demonstrating its reliability in noisy settings.
@@ 4.5.2 Resilience to Limited Training Data
Many medical datasets are limited in size due to privacy constraints. As shown in Figure 4(e), HDC-X remains robust under reduced training data, with only a 1.78% drop in accuracy when trained on 40% of the PhysioNet 2016 dataset.
@@ 4.5.3 Resilience to Hardware Errors
To evaluate this robustness in HDC-X, we conducted a perturbation analysis by randomly flipping elements (+1 to -1 and vice versa) in all stored Cluster-HVs to simulate hardware instability. Theorem 3 demonstrates that, with sufficient dimensionality, flipping up to 50% of the elements has minimal impact on classification accuracy (Proof 3). The theorem is supported empirically: as shown in Figure 4(f), with D=10, [^0] D=10{,}000, flipping 20% of elements results in only a 2.84% drop in accuracy.
Figure 4: HDC-X sensitivity to hyperparameters, input noise, limited training data, and hardware errors on PhysioNet 2016.
@@ 4.6 Conceptual Hardware Framework
Figure 5: Conceptual hardware framework for HDC-X.
The current implementation of HDC-X is in Python, which does not fully exploit the hardware-level efficiency of binary hypervectors enabled by single-bit operations and parallel computing. Prior work on HDC hardware has shown that fundamental operations such as binding and bundling can be implemented with extremely lightweight digital logic: binding reduces to bitwise XOR, bundling to majority voting circuits, and similarity search to parallelized Hamming distance computation (Imani et al., 2017). These properties allow HDC systems to operate with low energy consumption and high throughput compared to conventional floating-point ML models.
Building on these insights, we outline a conceptual hardware framework specifically tailored for HDC-X (Figure 5). To validate its feasibility, we prototyped the framework on a Xilinx FPGA platform and are able to achieve the accuracy reported in Table 1. We plan to further refine and evaluate this hardware-oriented design in future work.
@@ 5 Conclusion
This paper introduces HDC-X, an energy-efficient classification framework extended from Hyperdimensional Computing (HDC). HDC-X demonstrates significant advantage in medical data classification: on heart sound classification, it is $350\times$ more efficient than Bayesian ResNet, and provide $10\%$ accuracy improvement over standard HDC.
We evaluated HDC-X’s robustness both theoretically and empirically. The model maintains high accuracy under challenging conditions: showing only a 1.39% drop under 15% input noise, 1.78% when trained on just 40% of the dataset, and 2.84% when 20% of its learned parameters are corrupted. These results highlight HDC-X’s reliability on real-world medical applications.
From a societal perspective, HDC-X can help expand access to medical screening in underserved settings by enabling assessments on low-cost, GPU-free devices. However, overreliance on the screening tool without clinical oversight could lead to misdiagnoses; therefore, the system should be designed to recommend follow-up with a medical professional when abnormalities are detected.
@@
We are committed to ensuring the reproducibility of our results. To this end, we have released the complete HDC-X codebase along with a detailed installation guide at this GitHub repository. The repository contains all datasets used in our experiments as well as the exact hyperparameter configurations adopted for training and evaluation. This enables researchers to fully replicate our experiments and verify the reported results.